the sides of a triangle are in ratio 1:root3:2,then angles of a triangle are in ratio

Grade:12

1 Answers

Muhammad Kamran khan
4 years ago
Let  the sides of triangle be x
 Then a=x, b=√3x, c= 2x
We know that, a/sinA =b/sinB =c/sinC =x
Putting the value of a,b,c in above formula
x/sinA =√3x/sinB =2x/sinC =2x
1/sinA =√3/sinB =2/sinC =2
Now, 1/sinA =2,√3/sinB =2, 2/sinC =2
SinA=1/2, sinB=√3/2 sinC=1
A=30°,B=60°,C=90°
A:B:C=30°:60°:90°=1:2:3
It was easy